Please help solve
z=4/3(pi)r^3 find (pi)

Answer by stanbon(75887)

z=(4/3)(pi)r^3, find (pi)
Multiply both sides by (3/4) to get:
(3/4)z = pi*r^3
Divide both sides by r^3 to get:
pi = [3z/4r^3]
=================
Cheers,
Stan H.